# SSHPool: The Separated Subgraph-based Hierarchical Pooling

### v1

@@ III Methodology
To overcome the theoretical drawbacks of existing hierarchical pooling approaches, in this section we propose a novel Separated Subgraph-based Hierarchical Pooling (SSHPool) to learn effective graph representations. Similar to the other hierarchical pooling methods, our proposed approach can be applied for the GNN models. We first give the definition of the proposed SSHPool. Moreover, we develop an end-to-end GNN-based framework associated with the proposed SSHPool module for graph classification. Finally, we analyze the theoretical properties of the proposed SSHPool.
III-A The Separated Subgraph-based Hierarchical Pooling
@@
In this subsection, we propose a novel pooling method SSHPool, that is defined based on hierarchical separated subgraphs to reduce the over-smoothing problem arising in most existing hierarchical pooling methods. The overview of the proposed SSHPool for each computational layer is shown in Figure 2, and the computational architecture of the proposed SSHPool for each layer mainly consists of two steps. First, given an input graph, the nodes are adaptively assigned into different clusters represented by different colors. This process in turn generate a family of separated subgraphs, where each subgraph consists of the nodes belonging to the same cluster and remains the edge connection between these nodes from the original graph. Second, to extract the local structural information of each subgraph, we individually employ several Local Graph Convolution (LG Conv.) units to propagate the node information between each individual subgraph and then aggregate the subgraph into a coarsened node. Note that, the parameters for the local convolution operations of different subgraphs are not shared. Since each local convolution unit is restricted in each subgraph and the node information cannot be propagated between different subgraphs, significantly reducing the over-smoothing problem and providing discriminative coarsened node representations for the resulting coarsened graphs, i.e., the input graph for the next layer of SSHPool. By stacking the above computational steps (i.e., the coarsened graph will be employed as the new input graph for the next layer), the proposed SSHPool can hierarchically extract the intrinsic graph characteristics. Below, we give the detailed definitions for the proposed SSHPool of each layer.
Figure 2: The computational architecture for the proposed SSHPool of each layer.
@@
The Local Graph Convolution. For each layer $l$ of the proposed SSHPool, assume the input graph is denoted as $G^{(l)}(V^{(l)},E^{(l)})$ and has $n_{l}$ nodes in $V^{(l)}$ (i.e., $n_{l}=|V^{(l)}|$). Note that, $G^{l}$ can be either the original input graph (for $l=1$) or the coarsened graph from the last layer $l-1$ of the SSHPool (for $l\geq 2$). The first step of the SSHPool is to decompose $G^{(l)}$ into a family of $n_{l+1}$ separated subgraphs, by assigning the $n_{l}$ nodes of $G^{(l)}$ into $n_{l+1}$ clusters through an node assignment matrix. Unlike the classical hierarchical based DiffPool [17], the proposed SSHPool is based on the hard assignment matrix, i.e., each node cannot be assigned to multiple clusters. In fact, the hard assignment matrix can be computed through the soft one. Specifically, given the node feature matrix $X^{(l)}\in\mathbb{R}^{n_{l}\times d}$ of $G^{(l)}$, the soft assignment matrix $S_{\mathrm{soft}}$ is defined as
@@
where $S_{\mathrm{soft}}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l+1}}$ and $n_{l+1}<n_{l}$. With $S_{\mathrm{soft}}^{(l)}$ to hand, the $(i,j)$-th entry of the hard assignment matrix $S^{(l)}\in\{0,1\}^{n_{l}\times n_{l+1}}$ satisfies
@@
Clearly, each $i$-th row of the hard assignment matrix $S^{(l)}$ selects the maximum element as $1$ and sets the remaining elements as 00, i.e., the $i$-th node is only assigned to the $j$-th cluster. With $S^{(l)}$ to hand, we can decompose the input graph $G^{(l)}$ into $n_{l+1}$ separated subgraphs, and each $j^{\mathrm{th}}$ subgraph is denoted as $G_{j}^{(l)}(V_{j}^{(l)},E_{j}^{(l)})$, where the node set $V_{j}^{(l)}$ consists of the nodes belonging to the $j$-th cluster and the edge set $E_{j}^{(l)}$ remains the original edge connections between the nodes in $V_{j}^{(l)}$ from the original input graph $G^{(l)}$.
With the associated node feature matrix $X_{j}^{(l)}\in\mathbb{R}^{n_{l}\times d}$ and adjacent matrix $A_{j}^{(l)}\in\mathbb{R}^{|V_{j}^{(l)}|\times|V_{j}^{(l)}|}$ to hand, we propose an local graph convolution unit to extract the local structural information of each $j^{th}$ separated subgraph $G_{j}^{(l)}$ as
$Z^{(l)}_{j}=A_{j}^{(l)}X_{j}^{(l)}W_{j}^{(l)},$ (6)
where $W_{j}^{(l)}\in\mathbb{R}^{d\times d}$ is the trainable weight matrix of layer $l$, and $Z^{(l)}_{j}\in|V_{j}^{(l)}|\times d$ is the resulting local structural representation of $G_{j}^{(l)}$. Moreover, we define a mapping vector $\textbf{s}^{(l)}_{i}$ to compress each subgraph $G_{j}^{(l)}$ into a coarsened node, and
$\textbf{s}^{(l)}_{j}=\mathrm{softmax}(A_{j}^{(l)}X_{j}^{(l)}D_{j}^{(l)}),$ (7)
where $D_{j}^{(l)}\in\mathbb{R}^{d\times 1}$ is a trainable parameter vector, and will play a role to compress each $j^{\mathrm{th}}$ subgraph $G_{j}^{(l)}$ to the coarsened node in the following coarsened graph generation procedure for the next layer $l+1$ of the proposed SSHPool.
The Coarsened Graph Generation. After several individual local graph convolution operations on the separated subgraphs within the same layer of the proposed SSHPool, we aggregate the local information of different subgraphs to further generate a coarsened graph, as the input graph $G^{(l+1)}$ for the next layer of the proposed SSHPool. To this end, we collect the node feature matrices $Z^{(l)}_{j}$ of all the $j^{\mathrm{th}}$ subgraphs $G_{j}^{(l)}$ as the extracted node feature matrix $Z^{(l)}\in\mathbb{R}^{n_{l}\times d}$ for the original global graph $G^{(l)}$, where the node sequences of $Z^{(l)}$ follow that of the original graph $G^{(l)}$ and each vertex feature vector of $Z^{(l)}$ is equal to that of the corresponding node in $G_{j}^{(l)}$. This is because each node of $G_{j}^{(l)}$ essentially corresponds to an unique original node in $G^{(l)}$. With the hard assignment matrix $S^{(l)}$ defined by Eq.(5)and the node mapping vector $\textbf{s}^{(l)}_{i}$ of each $G_{j}^{(l)}$ defined by Eq.(7) to hand, the node feature matrix $X^{(l+1)}\in\mathbb{R}^{n_{l+1}\times d}$ and the adjacency matrix $A^{(l+1)}\in\mathbb{R}^{n_{l+1}\times n_{l+1}}$ of the resulting coarsened graph $G^{(l+1)}$ are defined as follows, i.e.,
$X^{(l+1)}=\mathrm{Reorder}[\mathop{\|}\limits_{j=1}^{n_{l+1}}\textbf{s}^{(l)^{T}}_{j}]Z^{(l)},$ (8)
$A^{(l+1)}=\alpha S^{(l)^{T}}A^{(l)}S^{(l)},$ (9)
@@
The Global Feature Generation. The above definitions give the computational procedures of the proposed SSHPool for each layer $l$. By hierarchically stacking the computational procedures with multiple layers, the initial graph $G^{(0)}$, the proposed SSHPool can hierarchically extract final global feature X for $G^{(0)}$ in the last layer $L$ ($l\leq L$). Specifically, given the initial input graph $G^{(0)}$ associated with its node feature matrix $X^{(0)}$ and the adjacent matrix $A^{(0)}$, the whole process of the proposed SSHPool can be generalized as
$\textit{{X}}=\mathrm{SSHPooling}_{L}(X^{(0)},A^{(0)};\Theta_{\mathrm{{S}}}),$ (10)
where $\mathrm{SSHPool}_{L}(\cdot)$ is the the proposed SSHPool function with $L$ layers, and $\Theta_{S}$ represents the parameter set of SSHPool.
@@ III-B The GNN Framework associated with the Proposed SSHPool
Based on the proposed SSHPool method, we develop an end-to-end GNN framework for graph classification. The computational architecture of our GNN framework is shown in Figure 3, and has three main computational steps. First, for each original input graph, we commence by performing a standard global graph convolution operation [8] to generate the initial node feature matrix as the graph embedding. Second, with the graph embedding of each original graph to hand, we perform the proposed SSHPool to further extract global hierarchical structural feature through multiple sets of layer-wise hierarchical separated subgraphs. Third, to overcome the possible degradation problem caused by the multiple hierarchical payers of the proposed SSHPool, we define an attention-based layer to further integrate the graph embedding and the global feature from the first and second steps, and the integrated graph feature is employed by the followed MLP for classification. Below we give the detailed definition of our attention-based layer.
Figure 3: The computational architecture of our GNN framework.
The Graph Attention Layer. After the hierarchical pooling process of the SSHPool operation, the original graph feature has been transformed into the global feature in a hierarchical structure. Since there are multiple layer-wise stacked local convolution unit sets associated with the hierarchical substructures of the proposed SSHPool (i.e., the SSHPool module may be deeper), resulting in possible degradation problems. To resolve this problem, we propose a graph attention layer to further integrate the initial graph embedding extracted from the global graph convolution operation and the global feature extracted from the proposed SSHPool. Given the global feature X defined by Eq.(10) and the original graph embedding $X^{(0)}$, the function of the attention-based mechanism is able to reinforce the key graph feature and learn the cross-information, and the corresponding attention operation is defined as
$\mathcal{A}=\mathrm{Attention}(X^{(0)},\textit{{X}})=\mathrm{softmax}(\frac{(\textit{{X}}W_{q})(X^{(0)}W_{k})^{T}}{\sqrt{d}})(X^{(0)}W_{v}),$ (11)
where $W_{k},W_{q},W_{v}\in\mathbb{R}^{d\times d}$ are the corresponding parameter matrices, and $\mathcal{A}$ is the attention value. With the graph attention layer to hand, the proposed SSHPool can also encapsulate the initial graph embedding from the original graph structure input to our GNN framework, so that the parameters related to the input graph maintain their values and avoid the gradient vanishing caused by the degradation. Finally, we adopt the MLP as the classifier. Given the attention matrix $\mathcal{A}$ defined by Eq.11 and the label set $\mathcal{Y}$, our GNN framework model computes the predicted graph labels $\hat{Pr}$ as
@@
where $\hat{Pr}=[\hat{p}_{1},\hat{p}_{2},...,\hat{p}_{\|\mathcal{Y}\|}]^{T}$ encapsulates the probabilities whose amount is $\|\mathcal{Y}\|$, and $\mathrm{Linear}(\cdot)$ is the active function of the MLP. In this work, we adopt the cross entropy as the loss function to train our GNN framework.
@@ III-C Discussions of the Proposed SSHPool
Compared to the aforementioned hierarchical-based graph pooling methods in Section I and Section II-B, the proposed SSHPool has some important properties, explaining the effectiveness. First, as we have stated in Section I and Section III-A, unlike the classical hierarchical-based DiffPool, ABDPool, ASAPool, the proposed SSHPool is defined on separated subgraphs, where there is no connection between the substructures. Thus, the proposed local graph convolution operation can only propagate the node structure information in each individual node cluster, i.e., the node information cannot be propagated between the nodes from different subgraphs. This can significantly reduce the over-smoothing problem and provide more discriminative structural characteristics for graphs. Since the proposed SSHPool focuses more on separated substructures, the resulting global graph features can encapsulate rich intrinsic structural information. Second, to overcome the possible degradation problem that maybe caused by the multiple layers of the proposed SSHPool, we also propose an attention layer for our GNN framework associated with the proposed SSHPool to simultaneously capture the initial graph embedding from the original graph structure, addressing the gradient vanishing caused by the degradation.
@@ IV Experiments
In this section, we evaluate the classification performance of our GNN-based framework associated with the proposed SSHPool on five standard graph datasets  [^1] http://graphlearning.io/. Details of these datasets are shown in Table I.
@@
Datasets D&D PROTEINS ENZYMES IMDB-B IMDB-M
Nodes(max.) 5748 620 126 136 89
Nodes(mean) 284.3 39.06 32.63 19.77 13.00
Graphs 1178 1113 600 1000 1500
Node labels 82 61 3 null null
Classes 2 2 6 2 3

### v2

@@ III The Methodology
To overcome the theoretical drawbacks of the existing hierarchical pooling approaches, in this section, we propose a novel Separated Subgraph-based Hierarchical Pooling (SSHPool) to learn effective graph representations. Similar to the other hierarchical pooling methods, our proposed approach can be applied to the GNN models. We first give the definition of the proposed SSHPool. Moreover, we develop an end-to-end GNN-based framework associated with the proposed SSHPool module for graph classification. Finally, we analyze the theoretical properties of the proposed SSHPool.
III-A Definitions of the Proposed SSHPool
@@
In this subsection, we propose a novel pooling method SSHPool, that is defined based on hierarchical separated subgraphs to reduce the over-smoothing problem arising in most existing hierarchical pooling methods. The overview of our proposed SSHPool for each computational layer is shown in Figure 2, and the computational architecture of the proposed SSHPool for each layer mainly consists of two steps. First, given an input graph, the nodes are adaptively assigned into different clusters represented by different colors. This process in turn generates a family of separated subgraphs, where each subgraph consists of the nodes belonging to the same cluster and retains the edge connection between these nodes from the original graph. Second, to extract the local structural information of each subgraph, we individually employ several Local Graph Convolution (LG Conv.) units to propagate the node information between each individual subgraph and then aggregate the subgraph into a coarsened node. Note that, the parameters for the local convolution operations of different subgraphs are not shared. Since each local convolution unit is restricted in each subgraph and the node information cannot be propagated between different subgraphs, significantly reducing the over-smoothing problem and providing discriminative coarsened node representations for the resulting coarsened graphs, i.e., the input graph for the next layer of SSHPool. By stacking the above computational steps (i.e., the coarsened graph will be employed as the new input graph for the next layer), the proposed SSHPool can hierarchically extract the intrinsic graph characteristics. Below, we give the detailed definitions for the proposed SSHPool of each layer.
Figure 2: The computational architecture for the proposed SSHPool of each layer. The first step is the assignment where the input graph is decomposed into multiple separated subgraphs. Then, the second one is the coarsening process where these subgraphs are compressed into the nodes of the coarsened graph.
@@
The Assignment and the Local Graph Convolution. For each layer $l$ of the proposed SSHPool, assume the input graph is denoted as $G^{(l)}=(V^{(l)},E^{(l)})$ and has $n_{l}$ nodes in $V^{(l)}$ (i.e., $n_{l}=|V^{(l)}|$). Note that, $G^{(l)}$ can be either the input graph embedded from the global convolution (for $l=1$) or the coarsened graph from the last layer $l-1$ of the SSHPool (for $l\geq 2$). The first step of the SSHPool is to decompose $G^{(l)}$ into a family of $n_{l+1}$ separated subgraphs, by assigning the $n_{l}$ nodes of $G^{(l)}$ into $n_{l+1}$ clusters through an node assignment matrix. And we define the assignment ratio as $\alpha=\frac{n_{l+1}}{n_{l}}$. Unlike the classical hierarchical based DiffPool [17], the proposed SSHPool is based on the hard assignment matrix, i.e., each node cannot be assigned to multiple clusters. In fact, the hard assignment matrix can be computed through the soft one. Specifically, given the node embedding matrix $X^{(l)}\in\mathbb{R}^{n_{l}\times d}$ of $G^{(l)}$, the soft assignment matrix $S_{\mathrm{soft}}^{(l)}$ is defined as
@@
where $S_{\mathrm{soft}}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l+1}}$ ($n_{l+1}<n_{l}$), and $\mathrm{softmax}(\cdot)$ is the softmax classifier. With $S_{\mathrm{soft}}^{(l)}$ to hand, the $(i,j)$-th entry of the hard assignment matrix $S^{(l)}\in\{0,1\}^{n_{l}\times n_{l+1}}$ satisfies
@@
Clearly, each $i$-th row of the hard assignment matrix $S^{(l)}$ selects the maximum element as $1$ and the remaining elements as 00, i.e., the $i$-th node is only assigned to the $j$-th cluster. With $S^{(l)}$ to hand, we can decompose the input graph $G^{(l)}$ into $n_{l+1}$ separated subgraphs, and each $j^{\mathrm{th}}$ subgraph is denoted as $G_{j}^{(l)}=(V_{j}^{(l)},E_{j}^{(l)})$, where the node set $V_{j}^{(l)}$ consists of the nodes belong to the $j$-th cluster and the edge set $E_{j}^{(l)}$ remains the original edge connections between the nodes in $V_{j}^{(l)}$ from the original input graph $G^{(l)}$.
With the associated node feature matrix $X_{j}^{(l)}\in\mathbb{R}^{|V_{j}^{(l)}|\times d}$ and the adjacent matrix $A_{j}^{(l)}\in\mathbb{R}^{|V_{j}^{(l)}|\times|V_{j}^{(l)}|}$ of each $j$-th separated subgraph $G_{j}^{(l)}$ to hand, we propose a local graph convolution unit to extract the local structural information as
$Z^{(l)}_{j}=\tilde{A}_{j}^{(l)}X_{j}^{(l)}W_{j}^{(l)},$ (6)
where $\tilde{A}_{j}^{(l)}=A_{j}^{(l)}+I$, $W_{j}^{(l)}\in\mathbb{R}^{d\times d}$ is the trainable weight matrix of layer $l$, and $Z^{(l)}_{j}\in\mathbb{R}^{|V_{j}^{(l)}|\times d}$ is the resulting local structural representation of $G_{j}^{(l)}$.
The Coarsened Graph Generation. After several individual local graph convolution operations on the separated subgraphs within the same layer of the proposed SSHPool, we aggregate the local information of different subgraphs to further generate a coarsened graph, as the input graph $G^{(l+1)}$ for the next layer of the proposed SSHPool. To this end, we collect the embedding local structural matrices of all the subgraphs, and compress each subgraph into the coarsened node.
In order to map the subgraphs to the coarsened graph nodes, we define a mapping vector $s^{(l)}_{j}$ to compress each subgraph $G_{j}^{(l)}$ into a coarsened node, and $s^{(l)}_{j}=S^{(l)}[:,j]$ where we split the hard assignment matrix into vector. And $s^{(l)}_{j}$ represents the nodes belonging to $j$-th subgraph. Note that, $s^{(l)}_{j}$ is a vector with the empty nodes which are not in the $j$-th subgraph. So we transfer $s^{(l)}_{j}$ into $\hat{s}^{(l)}_{j}$ where the empty nodes are removed and the rest of the nodes remain in order. With the hard assignment matrix $S^{(l)}$ defined by Eq.(5) and the mapping vector $\hat{s}^{(l)}_{j}$ of each $G_{j}^{(l)}$ to hand, the node feature matrix $X^{(l+1)}\in\mathbb{R}^{n_{l+1}\times d}$ and the adjacency matrix $A^{(l+1)}\in\mathbb{R}^{n_{l+1}\times n_{l+1}}$ of the resulting coarsened graph $G^{(l+1)}$ are defined as follows, i.e.,
$X^{(l+1)}=\mathop{\|}\limits_{j=1}^{n_{l+1}}\hat{s}^{(l)\top}_{j}Z^{(l)}_{j},$ (7)
$A^{(l+1)}=S^{(l)\top}A^{(l)}S^{(l)},$ (8)
@@
The Global Feature Generation. The above definitions give the computational procedures of the proposed SSHPool for each layer $l$. By hierarchically stacking the computational procedures with multiple layers, the initial graph $G^{(0)}$, the proposed SSHPool can hierarchically extract the final global feature X for $G^{(0)}$ in the last layer $L$ ($l\leq L$). Specifically, given the initial input graph $G^{(0)}$ associated with its node feature matrix $X^{(0)}$ and the adjacent matrix $A^{(0)}$, the whole process of the proposed SSHPool can be generalized as
$\textit{{X}}=\mathrm{SSHPool}_{L}(X^{(0)},A^{(0)};\Theta_{\mathrm{{S}}}),$ (9)
where $\mathrm{SSHPool}_{L}(\cdot)$ is the proposed SSHPool function with $L$ layers, and $\Theta_{S}$ represents the parameter set of SSHPool.
Figure 3: The computational architecture of our GNN framework.
@@ III-B The GNN Framework associated with the Proposed SSHPool
Based on the proposed SSHPool method, we develop an end-to-end GNN framework for the graph classification. The computational architecture of our GNN framework is shown in Figure 3, and has three main computational steps. First, for each original input graph, we commence by performing a standard global graph convolution operation [8] to generate the initial node feature matrix as the graph embedding. Second, with the graph embedding of each original graph to hand, we perform the proposed SSHPool to further extract the global hierarchical structural feature through multiple sets of layer-wise hierarchical separated subgraphs. Third, to overcome the possible degradation problem caused by the multiple hierarchical structure of the proposed SSHPool, we introduce an attention-based layer to further integrate the graph embedding and the global feature from the first and second steps, and the integrated graph feature is employed by the followed MLP for classification. Below we give the detailed definition of our attention-based layer.
The Graph Attention Layer. After the hierarchical pooling process of the SSHPool operation, the original graph feature has been transformed into the global feature in a hierarchical structure. Since there are multiple layer-wise stacked local convolution unit sets associated with the hierarchical substructures of the proposed SSHPool (i.e., the SSHPool module may be deeper), resulting in possible degradation problems. To resolve this problem, we propose a graph attention layer to further integrate the initial graph embedding extracted from the global graph convolution operation and the global feature extracted from the proposed SSHPool. Given the global feature X defined by Eq.(9) and the graph embedding $X^{(0)}$ from the global convolution operation, the function of the attention-based mechanism is able to reinforce the key graph feature and learn the cross-information, and attention calculation is defined as
$\mathcal{A}$ $=\mathrm{Attention}(X^{(0)},\textit{{X}})$ (10)
$=\mathrm{softmax}\left(\frac{(\textit{{X}}W_{q})(X^{(0)}W_{k})^{\top}}{\sqrt{d}}\right)(X^{(0)}W_{v}),$
where $W_{k},W_{q},W_{v}\in\mathbb{R}^{d\times d}$ are the corresponding parameter matrices, and $\mathcal{A}$ is the attention value. With the graph attention layer to hand, the proposed SSHPool can also encapsulate the initial graph embedding from the original graph structure input to our GNN framework, so that the parameters related to the input graph maintain their values and avoid the gradient vanishing caused by the degradation. Finally, we adopt the MLP as the classifier. Given the attention matrix $\mathcal{A}$ defined by Eq.(10) and the label set $\mathcal{Y}$, our GNN framework model computes the predicted graph labels as
@@
where $\hat{Pr}=[\hat{p}_{1},\hat{p}_{2},...,\hat{p}_{\|\mathcal{Y}\|}]^{T}$ encapsulates the probabilities whose amount is $\|\mathcal{Y}\|$, and $\mathrm{Linear}(\cdot)$ is the active function of the MLP. And we adopt the cross entropy as the loss function in our training process.
@@ III-C Discussions of the Proposed SSHPool
Compared to the aforementioned hierarchical-based graph pooling methods in Section I-A and Section II-B, the proposed SSHPool has some important properties, explaining the effectiveness. First, as we have stated in Section I-A and Section III-A, unlike the classical hierarchical-based DiffPool, ABDPool, ASAPool, the proposed SSHPool is defined on the separated subgraphs, where there is no connection between the substructures. Thus, the proposed local graph convolution operation can only propagate the node structure information in each individual node cluster, i.e., the node information cannot be propagated between the nodes from different subgraphs. This can significantly reduce the over-smoothing problem and provide more discriminative structural characteristics for graphs. Since the proposed SSHPool focuses more on separated substructures, the resulting global graph features can encapsulating rich intrinsic structural information. Second, to overcome the possible degradation problem that maybe caused by the multiple layers of the proposed SSHPool, we also propose an attention layer for our GNN framework associated with the proposed SSHPool to simultaneously capture the initial graph embedding from the original graph structure, addressing the gradient vanishing caused by the degradation.
@@ IV Experiments
In this section, we evaluate the classification performance of our GNN-based framework associated with the proposed SSHPool on seven standard graph datasets  [^1] http://graphlearning.io/. Details of these datasets are shown in Table I.
@@
Datasets Vertices(max.) Vertices(mean.) Graphs Vertex labels Classes
D&D 5748 284.32 1178 82 2
PTC 64 14.29 344 18 2
PROTEINS 620 39.06 1113 3 2
NCI1 111 29.87 4110 37 2
NCI109 111 29.68 4127 38 2
IMDB-B 136 19.77 1000 null 2
IMDB-M 89 13.00 1500 null 3